Your company has 20 million shares outstanding, total earnings this year of $50 million, and a 20% payout ratio.

a. If your return on new investment is 11% and you maintain your payout ratio at 20%, what will be next year’s dividend per share? [7]

b. Now assume that you reduce your payout ratio so that this year’s payout rate will be 10%, and next year’s payout ratio will also be 10%. At what rate will your dividends grow? Compute both dividends and compare them to calculate the growth rate.
Answer #1

Growth Rate = Retention Ratio * RoE

= (100% - 20% ) * 11%

(G) = 8.8%

Earnings next year = 50 Mn ( 1+G)

= 50 ( 1+ 0.088)

= 54.4 Mn

Dividends = Earnings * Payout Ratio

= 54.4 Mn * 20%

= 10.88 Mn

Dividend per share = 10.88 Mn / 20 Mn

   = 0.544 per share  

Answer b)

Growth Rate = Retention Ratio * RoE

= (100% - 10% ) * 11%

(G) = 9.9%

Earnings next year = 50 Mn ( 1+G)

= 50 ( 1+ 0.099)

= 54.95 Mn

Dividend this year = (Earnings this year * Payout Ratio)/ Outstanding shares

= (50 Mn * 10%) / 20

= $ 0.25

Dividend next year = (Earnings next year * Payout Ratio)/ Outstanding shares

= (54.95 Mn * 10%) / 20

= $ 0.27475

Growth Rate of Dividend =  (Dividend next year - Dividend this year) / Dividend this year

= (0.27475 - 0.25) / 0.25

= 9.9%